OpenStudy (anonymous):

Can anyone help me divide a rational expression? Find the Quotient and Simplify. 7x^7 over 3x^5 divided by 14x over 9x^3

OpenStudy (phi):

if you divide by a fraction, you can "flip it" and multiply by it. Can you do that for this problem ?

OpenStudy (anonymous):

Yes, I know I have to flip the second fraction. I'm just not sure how to find the common denominator

OpenStudy (phi):

you don't need a common denominator when multiplying (or dividing) you should get \[ \frac{7x^7}{3x^5} \cdot \frac{9x^3}{14x} \] when multiplying you multiply top times top and bottom times bottom but you can first divide (it's usually easier)

OpenStudy (phi):

order can be changed when multiplying for example, the top can be written as 7*9 * x^7 * x^3 and at the bottom 14*3*x^5*x \[ \frac{ 7 \cdot 9 \cdot x^7 \cdot x^3}{14 \cdot 3 \cdot x^5 \cdot x}\] and you can divide the numbers

OpenStudy (phi):

you can "split apart" that fraction, and write it like this \[ \frac{7}{14} \cdot \frac{9}{3} \cdot \frac{x^7 \cdot x^3}{x^5\cdot x} \]

OpenStudy (phi):

people don't usually bother, but it might make it easier to see how to simplify the numbers can you simplify 7/14 ?

OpenStudy (anonymous):

yes that would be 1/2

OpenStudy (phi):

For the x's you use the idea that \[ \frac{x}{x} = 1\] anything divided by itself is 1 so if you see x^7 (which is short for x*x*x*x*x*x*x) and you see x^5 in the bottom: x*x*x*x*x you can see 5 of the x's in the bottom can divide into 5 in the top (you get 1*1*1*1*1) leaving x*x up top yes 7/14 is ½ what about 9/3 ?

OpenStudy (anonymous):

9/3 would be \[\frac{ 1 }{ 3 }\]

OpenStudy (phi):

other way round , right ?

OpenStudy (anonymous):

oops! yes, other way around

OpenStudy (phi):

ok, make a note: the numbers simplify to ½ * 3 or 3/2 can you do the x's ?

OpenStudy (anonymous):

I think so. Would the final answer be 3x^11 over 2^6?

OpenStudy (phi):

how did you get that ?

OpenStudy (phi):

ignore the numbers and look just at \[ \frac{x^7 x^3}{x^5x} \]

OpenStudy (anonymous):

Don't you add the exponents?

OpenStudy (phi):

with all those x's it's a pain to type out, so let's do a small problem \[ \frac{x^3}{x} \] as you know x^3 means x*x*x \[ \frac{x\cdot x\cdot x}{x} \] and as you know we can make that into separate fractions \[ \frac{x}{x} \cdot \frac{x}{1}\cdot \frac{x}{1} \] notice you multiply top times top and bottom times bottom we get what we started with

OpenStudy (phi):

we can simply by know x/x is 1 and 1 times anything is the anything in other words we get \[ \frac{x}{1}\cdot \frac{x}{1} = \frac{x^2}{1}= x^2 \] the fast way is to subtract the bottom exponent from the top exponent 3-1= 2 and write x^2 right away

OpenStudy (phi):

yes, you can add exponent when multiplying x^2 * x^2 means x*x * x*x or in short hand x^4 the fast way is to do 2+2 to get 4 for the exponent. so you could do the top and get x^(7+3) = x^10 and the bottom you get x^(5+1) = x^6 and you can divide x^6 into x^10 to get x^4

OpenStudy (anonymous):

Ok I'm getting a little lost now

OpenStudy (phi):

try this one: \[ \frac{x}{x} \] what does the simplify to ?

OpenStudy (anonymous):

1

OpenStudy (phi):

and you might know if we put in an exponent of 1 x^1 = x (we don't usually bother to show an exponent if it's one) \[ \frac{x^1}{x^1} \] if we subtract the exponents top - bottom : 1- 1=0 we get \[ \frac{x^1}{x^1} =x^0 =1\]

OpenStudy (anonymous):

I'm forgetting what my original problem is now

OpenStudy (phi):

anything to the 0 power is 1. (make a note of that... it may not be obvious)

OpenStudy (phi):

we will do you problem in just a minute can you simplify \[ \frac{x\cdot x}{x} \] ?

OpenStudy (phi):

notice it's the same as simplifying \[ \frac{x}{x} \cdot \frac{x}{1} \]

OpenStudy (anonymous):

it would still be 1

OpenStudy (anonymous):

or x^2 over 1x

OpenStudy (phi):

x/x is 1 you can "replace" the x/x with 1, but you would then have \[ \frac{x}{x} \cdot \frac{x}{1} =1 \cdot \frac{x}{1} \]

OpenStudy (phi):

This is the idea you want to learn \[ \frac{x\cdot x}{x} = \frac{x}{x} \cdot \frac{x}{1} = 1 \cdot \frac{x}{1} =\frac{x}{1}\]

OpenStudy (anonymous):

I'm running out of time and just need the answer or at least simple steps to the answer

OpenStudy (phi):

add the exponents in the top what do you get ?

OpenStudy (anonymous):

10

OpenStudy (phi):

now do the same in the bottom

OpenStudy (anonymous):

6

OpenStudy (phi):

and because you are dividing do 10 - 6 to get the exponent for x

OpenStudy (anonymous):

4

OpenStudy (phi):

in other words x^4 with the numbers you get \[ \frac{3x^4}{2} \]

OpenStudy (anonymous):

and that is found by multiplying the fractions and then simplifying them? Thank you for all your help

OpenStudy (mathstudent55):

You can also reduce each fraction before dividing them. \(\Large \dfrac{7x^7}{3x^5} \div \dfrac{14x}{9x^3} \) \(\Large =\dfrac{7x^2}{3} \div \dfrac{14}{9x^2} \) \(\Large =\dfrac{7x^2}{3} \times \dfrac{9x^2}{14} \) \(\Large =\dfrac{7 \times 3 \times 3 x^4}{3 \times 2 \times 7} \) \(\Large =\dfrac{\cancel{7} \times \cancel{3} \times 3 x^4}{\cancel{3} \times 2 \times \cancel{7}} \) \(\Large =\dfrac{3 x^4}{ 2}\)